OpenStudy (anonymous):

What are the factors of 3x2 + 10x + 8?

myininaya (myininaya):

ask yourself what are two factors of 3(8) that have product 3(8) and have sum 10

myininaya (myininaya):

3(8)=6(4) 10=6+4 10x=6x+4x replace 10x with 6x+4x and then factor by grouping